Find the electrostatic force between charges of -2.0 C and +5.0C separated by a distance of 75m in a vacuum.

1
Expert's answer
2022-02-17T10:05:26-0500

Given:

"q_1=-2.0\:\rm C"

"q_2=5.0\:\rm C"

"r=75\:\rm m"


The Coulomb's law says

"F=k\frac{|q_1||q_2|}{r^2}"

"F=9*10^9\frac{2.0*5.0}{75^2}=1.6*10^7 \:\rm N"

The force is attraction.